OpenStudy (anonymous):

which proprtion can be used to determine what number is 15% of 30? 15/30 = x/100 15/100= x/30 15/100=30/x 30/100=15/x

OpenStudy (lieutenantgeneral):

15/100 = 30/x 30/x = 30/15 = 3/20

OpenStudy (anonymous):

so which on is it